OpenStudy (campayne):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) {y = x - 19 2x - y = 27

OpenStudy (sushi121212):

If you plug in the y=x-19 into the y part of 2x-y=27, you will get the answer: 2x-(x-19)=27

OpenStudy (campayne):

which boils down to x = 8 ive been getting

OpenStudy (sushi121212):

Yes, then plug 8 back into one of the equations: y=x-19 y=8-19 y=-11 So it should be (8, -11) You can always plug if back in to check: 2x - y = 27 2(8)-(-11) = 27 16+11= 27 27 = 27 So the answer is (8, -11)

OpenStudy (campayne):

i thought so lol thats what i got